# Wave Function Collapse, Lorentz Invariance, and the Third Postulate of Relativity

### v1

@@ III Interaction-Induced Wave Function Collapse
The adoption of this sort of additional stochastic structure transfers the probabilistic nature of quantum theory from the macroscopic to the elementary level. This turns what were ad hoc postulates at the macro level into straightforward consequences of the mathematical description of elementary processes. The probabilistic nature of elementary correlating interactions also explains why a preferred reference frame (or spacetime foliation) remains hidden, and explains why the description of spacetime at the macroscopic level is limited to the standard relativistic account involving just a Lorentzian metric. This is why accounts involving these sorts of additional stochastic structures ought to be regarded as respecting serious Lorentz invariance.

### v2

@@ IV Reconciling Relativity and Quantum Nonlocality
The adoption of this sort of additional stochastic structure transfers the probabilistic nature of quantum theory from the macroscopic to the elementary level. This turns what were ad hoc postulates at the macro level into straightforward consequences of the mathematical description of elementary processes. The probabilistic nature of elementary correlating interactions also explains why a preferred reference frame (or spacetime foliation) remains hidden, and explains why the description of spacetime at the macroscopic level is limited to the standard relativistic account involving just a Lorentzian metric. This is why accounts involving these sorts of additional stochastic structures ought to be regarded as genuinely relativistic.
@@
Acknowledgements: I would like to thank Nicolas Gisin and Flavio Del Santo for useful discussions.